OpenStudy (anonymous):

For the system shown below, what are the coordinates of the solution that lies in quadrant II? x^2+4y^2=100 4y-x^2=-20 PLEASE HELP FAST!!!!!

OpenStudy (campbell_st):

well equate them then solve \[x^2 + 4y^2 = 100\] and \[x^2 = 4y + 20\] so substituting the 2nd equation into the 1st you get \[4y + 20 + 4y^2 = 100\] which can be written as \[4y^2 + 4y - 80 = 0\] divide each term by 4 and you get \[y^2 + y - 20 =0\] now solve for y

OpenStudy (anonymous):

(y+5)(y-4)

OpenStudy (anonymous):

so y is -5 or 4

OpenStudy (anonymous):

Now what??? @campbell_st????

OpenStudy (campbell_st):

ok... so for the 2nd quadrant... you'll need to positive y so substitute it into either equation, and use the negative x answer for the ordered pair

OpenStudy (anonymous):

thank you!!!!